OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (math&ing001):

Add (-2) to the first term and go on like this -10, -10-2=-12, -12-2=-14.........

OpenStudy (anonymous):

thanks you @math&ing001

OpenStudy (math&ing001):

you're welcome